# Machine Space I: Weak exponentials and quantification over compact spaces

### v4

@@ Abstract.
Key words and phrases:
hyperspace, continuous lattice, classifying locale
2020 Mathematics Subject Classification:
54B20, 54-08, 54C35, 06D22, 06B35, 54D30
@@ 1. Introduction
In summary, the logic of verifiability admits finite conjunctions and arbitrary disjunctions. Identifying each proposition $U$ with the set of points which satisfy it recovers the usual notion of a topological space with conjunctions interpreted as set intersection and disjunctions as set union. Thus, point-set topology provides one formalisation of the theory of verifiability. Indeed, this perspective even extends to continuous functions.
@@
Concrete topological spaces can be fruitfully interpreted from the perspective of verifiability.
Example 1.1.
Consider the case of the real numbers $\mathbb{R}$. We might imagine having devices that can measure a quantity $x$, each having some fixed precision $\varepsilon>0$ and giving rational outputs. More precisely, for a real number $x$, such a device will (nondeterministically) give some $q\in\mathbb{Q}$ such that $\lvert x-q\rvert<\varepsilon$. Moreover, let us assume that for any precision there exists a device that is at least that precise.
@@
Given this perspective, it is possible to derive interpretations of various other topological concepts in terms of verifiability. For instance, it is well known that a compact set $K$ corresponds to a set that can be universally quantified over — that is, given a verifiable property $P$ it is possible to verify that each member of $K$ satisfies $P$ (Note that $K$ may in fact be infinite!)
@@ 2.1. Pointfree topology
As mentioned before, compactness has a very nice interpretation in terms of universal quantification of verifiable properties. We will explore this topic further in Section˜4.
@@ 2.2. The Scott topology and exponentials
The set of opens $\mathcal{O}X$ of a locale $X$ is itself often endowed with the Scott topology.
Definition 2.5.
A subset $V\subseteq\mathcal{O}X$ is Scott-open if it is upward closed and if whenever a join $\bigvee D$ of a directed set $D$ is in $V$, then some $d\in D$ lies in $V$. Recall that a set $D$ is directed if every finite subset of $D$ has an upper bound in $D$. ∎
The Scott topology can actually be defined more generally on dcpos, which are often used to model data types in programming languages (as discussed in Section˜4.3).
Definition 2.6.
A poset which admits joins of all directed subsets is called a dcpo. We will write $\operatorname*{\bigvee{}^{\!\uparrow}_{\hskip-0.60275pt\hskip 2.2604pt}}S$ for the join of a directed set $S$. A morphism of dcpos or a Scott-continuous function is a monotone map which preserves directed suprema. ∎
@@
Recall if $X,B$ are objects in category that the exponential object $B^{X}$ is equipped with a morphism $\mathrm{ev}\colon B^{X}\times X\to B$ such that for every $f\colon Y\times X\to B$ there is a unique $\overline{f}\colon Y\times X\to B^{X}$ such that the following diagram commutes.
@@
If $X$ is a locally compact locale, then $X$ is exponentiable (see [13]). In particular, the exponential locale $\Sigma^{X}$ exists and coincides with the Scott topology on $\mathcal{O}X\cong\mathrm{Hom}(X,\Sigma)$. Furthermore, $\Sigma^{X}$ is itself locally compact (for instance, see [28, Theorem 11.1]). However, if $X$ is not locally compact, the evaluation map $\mathrm{ev}\colon\mathcal{O}X\times X\to\Sigma$, which sends $(f,x)$ to $f(x)$, is not continuous when $\mathcal{O}X$ is given the Scott topology. In fact, in this case the exponential $\Sigma^{X}$ does not exist.
@@ 3.1. Presentations and weak exponentials
Recall that $\langle G\mid R\rangle$ is a quotient of the free frame on $G$, which in turn is the frame of opens of $\Sigma^{G}$. The frame quotient $\Sigma^{G}\twoheadrightarrow\langle G\mid R\rangle$ exhibits $X$ as a sublocale of $\Sigma^{G}$. The points of $\Sigma^{G}$ correspond to subsets of $G$ and we can then interpret this inclusion as sending points of $X$ to the set of generating opens in which they lie. The points in the image of this inclusion correspond to the frame homomorphisms $\mathcal{O}(\Sigma^{G})\to\mathcal{O}1$ which respect the relations. The additional points in $\Sigma^{G}$ may be thought of as ‘imaginary points’ in the same sense that the exponential spaces used in (the different approach of) [7, Chapter 10] are described as imaginary exponentials. Just as complex numbers may used to more easily prove results about real numbers, so will these imaginary points be useful for our study of compactness in Section˜4.
@@
Observe that $u$ defines an open in $A\times X$. Since $A\times X$ is a subspace of $A\times\Sigma^{G}$, there exists a (not necessarily unique) open $u^{\prime}$ in $A\times\Sigma^{G}$ which restricts to give $u$. By the universal property of the exponential $\Sigma^{\Sigma^{G}}$, we obtain a morphism $\overline{u}^{\prime}$ making the bottom right triangle commute. The left-hand trapezium commutes since both composites give $\overline{u}^{\prime}\times i_{X}$ and then a simple diagram chase confirms that $u=\widetilde{\mathrm{ev}}\circ(\overline{u}^{\prime}\times X)$. ∎
The above construction is summarised in the following definition.
Definition 3.3.
Let $X$ be a locale with $\mathcal{O}X=\langle G\mid R\rangle$. We define the associated (idealised) machine space to be $\Sigma^{\Sigma^{G}}$ together with the evaluation map $\widetilde{\mathrm{ev}}\colon\Sigma^{\Sigma^{G}}\times X\to\Sigma$ defined in Proposition˜3.2. ∎
We explore the intuition behind the concept of machine space in the following section.
@@ 3.2. Philosophical Interlude
Given a space $X$ with $\mathcal{O}X=\langle G\mid R\rangle$ we interpret each generator $g\in G$ as a basic black-box machine which takes as input the points of $X$ and either halts after some finite amount of time, or runs forever. More complicated machines can be constructed from these basic machines. For instance, we may construct a composite machine by taking a finite collection of machines and insist that each machine halts on the given point before the composite machine is considered to have halted. We can also construct a further machine by running any number of these composite machines in parallel and halting on a point if at least one branch halts. This corresponds to taking formal joins of finite formal meets of generators and so we see that these machines are precisely the elements of the frame frame on $G$ – that is, the points of the machine space $\Sigma^{\Sigma^{G}}$.
Definition 3.4.
@@
Notice that $\Sigma^{\Sigma^{G}}$ by itself contains no information about the behaviour of the machines on inputs from the locale $X$ with presentation $\langle G\mid R\rangle$, as this data is contained in the relation $R$. We can recover this information from the evaluation map $\widetilde{\mathrm{ev}}\colon\Sigma^{\Sigma^{G}}\times X\to\Sigma$ of Proposition˜3.2, which records whether a machine halts on a given point.
Remark 3.5.
In fact, for many purposes any weak exponential with base $\Sigma$ exponent $X$ behaves like a machine space for $X$. This could be interesting to explore further. However, for this paper we restrict our consideration to $\Sigma^{\Sigma^{G}}$, since the compactness algorithm given in Section˜4 is dependent on the presentation of the locale by generators and relations.
A point $x\in X$ is accepted by a machine $m$ if there exists some $J_{k}$ such that $x$ is accepted by each ‘basic machine’ $g_{j}$ for all $j\in J_{k}$ (i.e. such that $x$ lies in the open given by $g_{j}$ for all such $j$). In this way, each machine $m$ will define an open on which it halts. Many of these machines will correspond to the same opens, as described by the relation $R$, so that each open in $X$ corresponds to an equivalence class of machines. In Definition˜3.9 of Section˜3.3 we will see that when the exponential $\Sigma^{X}$ exists there is a quotient map $\Sigma^{\Sigma^{G}}\twoheadrightarrow\Sigma^{X}$ assigning machines to opens.
Remark 3.6.
@@
Recall that the failure of the exponential $\Sigma^{X}$ to always exist poses a puzzle for the verifiable interpretation of topology. Luckily, we have seen that machine spaces always exists. It is now helpful to consider in detail how such spaces might correspond to concrete objects that can be learnt about via physical measurements. Note that with such “real-world” models the set of generators is assumed to be countable.
Model 3.7.
@@
Note that associativity and distributivity are automatically satisfied with distributivity following from the fact that the formal operations are already reduced to the canonical form of a join of meets. These axioms together are then enough to give a frame ∎
@@
We can now answer the question posed in the introduction. While we have seen that machine space is reasonably concrete, the ‘space of opens’ is less so. In the flawed argument for the existence of a space of opens from Section˜1, one assumes that given an open and a point, it is possible to semi-decide membership of the point in the open. But an open merely describes the verifiable property under consideration — it is a machine that actually carries out this verification.
In the following subsection, we will show that the quotient map $\Sigma^{\Sigma^{G}}\twoheadrightarrow\mathcal{O}X$ has a section whenever the space $X$ is locally compact. From the above perspective, this fact perfectly explains why in case a continuous evaluation map exists for $\mathcal{O}X$: given an open $u$ and a point $x$, one can apply the section to $u$ to acquire a machine $s(u)$ which can then verify membership of $x$ in $u$.
@@
Remark 3.8.
Note that since we not care which specific machine representing a given open is selected, we can relax the requirement of having an actual section to having a sufficiently well-behaved quotient. In particular, suplattice and preframe homomorphisms between frames are known to correspond to angelically and demonically nondeterministic maps respectively [24, 31, 30] and it would seem that a nondeterministic choice of representative for each open would be sufficient to construct the evaluation map. Indeed, open and proper quotients, or more generally, triquotients, have such ‘non-deterministic sections’ and these are strong enough conditions on the quotient map to entail that $X$ is exponentiable. We will discuss the role of nondeterminism in more detail in a later paper.
@@ 3.3. Relationship between machine space and the space of opens
$\Sigma^{\Sigma^{G}}\times X$$\Sigma^{X}\times X$$\Sigma$$\Sigma^{X}$$\Sigma^{\Sigma^{G}}$$q\times X$$\mathrm{ev}$$\widetilde{\mathrm{ev}}$$q$
@@
Proposition 3.10.
Let $X$ be a locally compact locale with $\mathcal{O}X=\langle G\mid R\rangle$. The map $q\colon\Sigma^{\Sigma^{G}}\to\Sigma^{X}$ as defined in Definition˜3.9 has a section $s$ satisfying $\widetilde{\mathrm{ev}}\circ(s\times X)=\mathrm{ev}$.
Proof.
@@
By the diagram we have $\mathrm{ev}\circ(q\times X)\circ(s\times X)=\widetilde{\mathrm{ev}}\circ(s\times X)=\mathrm{ev}=\mathrm{ev}\circ(\mathrm{id}_{\Sigma^{X}}\times X)$. Hence by the uniqueness condition of the universal property of the ‘true’ exponential $\Sigma^{X}$, we have $q\circ s=\mathrm{id}_{\Sigma^{X}}$, as required. ∎
@@
Recall that if a locale $X$ is given by $\mathcal{O}X=\langle G\mid R\rangle$ then the frame $\mathcal{O}X$ is a quotient of $\mathcal{O}(\Sigma^{G})$ by the congruence generated from the relations $R$. Let us call this frame quotient map $\overline{q}$ for notational convenience, but as a function it is the same as $i_{X}^{*}$. Recall that $\mathcal{O}(\Sigma^{G})$ equipped with the Scott topology is precisely the machine space $\Sigma^{\Sigma^{G}}$. We can now use $\overline{q}$ to induce a topology on $\mathcal{O}X$. We will make use of the following result.
Lemma 3.11.
Let $L$ and $M$ be posets admitting all joins and let $q\colon L\twoheadrightarrow M$ be join-preserving surjection. Then if $L$ is equipped with the Scott-topology, the quotient topology induced on $M$ is also the Scott-topology.
Proof.
We need to show that $U\subseteq M$ is Scott-open if and only if $q^{-1}(U)$ is Scott-open.
Now suppose $q^{-1}(U)$ is Scott-open. Consider $a\in U$ and suppose $a\leq b$. Take $a^{\prime}\in q^{-1}(\{a\})$ and $b^{\prime}\in q^{-1}(\{b\})$. We may assume $b^{\prime}\geq a^{\prime}$ by joining $a^{\prime}$ if necessary. Then $a^{\prime}\in q^{-1}(U)$ and hence $b^{\prime}\in q^{-1}(U)$, since $q^{-1}(U)$ is upward closed. Thus, $b=q(b^{\prime})\in U$ and so $U$ is an upset. Next suppose $\operatorname*{\bigvee{}^{\!\uparrow}_{\hskip-0.60275pt\hskip 2.2604pt}}D\in U$. Set $D^{\prime}=q^{-1}({\downarrow}D)$. We note $D^{\prime}$ is directed. Certainly $D^{\prime}$ is nonempty. If $c,d\in D^{\prime}$ then there is an $e\in D$ with $q(c),q(d)\leq e$ by directedness of $D$. Thus, $q(c\vee d)=q(c)\vee q(d)\leq e\in{\downarrow}D$ and so $c\vee d\in D^{\prime}$. Furthermore, note that $q(D^{\prime})={\downarrow}D$ and $q(\operatorname*{\bigvee{}^{\!\uparrow}_{\hskip-0.60275pt\hskip 2.2604pt}}D^{\prime})=\operatorname*{\bigvee{}^{\!\uparrow}_{\hskip-0.60275pt\hskip 2.2604pt}}q(D^{\prime})=\operatorname*{\bigvee{}^{\!\uparrow}_{\hskip-0.60275pt\hskip 2.2604pt}}{\downarrow}D=\operatorname*{\bigvee{}^{\!\uparrow}_{\hskip-0.60275pt\hskip 2.2604pt}}D\in U$. Thus, $\operatorname*{\bigvee{}^{\!\uparrow}_{\hskip-0.60275pt\hskip 2.2604pt}}D^{\prime}\in q^{-1}(U)$ and so there is a $d^{\prime}\in D^{\prime}$ with $d^{\prime}\in q^{-1}(U)$, since $q^{-1}(U)$ is Scott-open. Now $q(d^{\prime})\in{\downarrow}D$ so that $q(d^{\prime})\leq d$ for some $d\in D$. Recalling that $U$ is an upset, we have $d\geq q(d^{\prime})\in U$ and so $d$ is the desired $d\in U$. Thus, $U$ is Scott-open. ∎
Corollary 3.12.
@@
Lemma 3.13.
The map $\overline{q}\colon\Sigma^{\Sigma^{G}}\to\mathcal{O}X$ is equal to $q\colon\Sigma^{\Sigma^{G}}\to\Sigma^{X}$ as defined in Definition˜3.9 whenever $X$ is locally compact.
Proof.
@@
The map $\mathrm{Hom}(i_{X},\Sigma)$ corresponds to the action of the frame map $i_{X}^{*}=\overline{q}$, while $\mathrm{Hom}(1,\Sigma^{i_{X}})$ gives the action of $\Sigma^{i_{X}}$ on points. Thus, we have that $\Sigma^{i_{X}}$ and $\overline{q}$ agree on points and hence they coincide. ∎
The failure of $\mathcal{O}X$ with the Scott topology to be an exponential $\Sigma^{X}$ when $X$ is not locally compact is due to this quotient map being badly behaved in general. Indeed, if there were a section to $q$ as in Proposition˜3.10 we could show $X$ to be locally compact after all.
Proposition 3.14.
@@
Proof.
Since $\Sigma^{G}$ is locally compact, $\mathcal{O}(\Sigma^{G})$ is a continuous frame and in particular a continuous dcpo. But continuous dcpos are known to be stable under Scott-continuous retracts and so $\mathcal{O}X$ is also a continuous dcpo (see [2, Theorem 3.1.4.]). Thus, $\mathcal{O}X$ is a continuous frame and $X$ is locally compact.
To see that $\widetilde{\mathrm{ev}}\circ(s\times X)$ is the evaluation map, first note that, by Lemma˜3.13, $\overline{q}=q$ as defined in Definition˜3.9. So by the commutative diagram in Definition˜3.9, the evaluation map $\mathrm{ev}\colon\mathcal{O}X\times X\to\Sigma$ satisfies $\mathrm{ev}\circ(q\times X)=\widetilde{\mathrm{ev}}$. Hence, $\widetilde{\mathrm{ev}}\circ(s\times X)=\mathrm{ev}\circ(q\times X)\circ(s\times X)=\mathrm{ev}$. ∎
In fact, we can say more (see Remark˜3.8). It is known that the lack of exponential objects is related to the fact that quotients are not stable under products (see [6]). We have the following result. (This result requires us to assume $X$ is spatial and that we are in the category of topological spaces, though it is still useful for intuition. But see also Remark˜3.16.)
Proposition 3.15.
@@
If $h\colon A\times X\to\Sigma$ then by the weak universal property of $\widetilde{\mathrm{ev}}$ we have a map $h^{\prime}\colon A\to\Sigma^{\Sigma^{G}}$ making the diagram commute. Then $h^{\prime\prime}=\overline{q}h^{\prime}$ satisfies the condition needed for the weak universal property of $\mathrm{ev}\colon\mathcal{O}X\times X\to\Sigma$ to hold. Moreover, it is unique since we have uniqueness on the underlying sets by the universal property of the exponential in $\mathrm{Set}$. Thus, $\mathcal{O}X$ satisfies the universal property of the exponential $\Sigma^{X}$, as required. ∎
Remark 3.16.

### v5

@@ 1. Introduction
In summary, the logic of verifiability admits finite conjunctions and arbitrary disjunctions. Identifying each proposition $U$ with the set of points which satisfy it recovers the usual notion of a topological space with verifiable properties corresponding to open sets, conjunctions interpreted as set intersection and disjunctions as set union. Thus, point-set topology provides one formalisation of the theory of verifiability. Indeed, this perspective even extends to continuous functions.
@@
Concrete topological spaces can be fruitfully interpreted from the perspective of verifiability. {exa} Consider the case of the real numbers $\mathbb{R}$. We might imagine having devices that can measure a quantity $x$, each having some fixed precision $\varepsilon>0$ and giving rational outputs. More precisely, for a real number $x$, such a device will (nondeterministically) give some $q\in\mathbb{Q}$ such that $\lvert x-q\rvert<\varepsilon$. Moreover, let us assume that for any precision there exists a device that is at least that precise.
@@
Given this perspective, it is possible to derive interpretations of various other topological concepts in terms of verifiability. For instance, it is well known that a compact set $K$ corresponds to a set that can be universally quantified over — that is, given a verifiable property $P$ it is possible to verify the property that all members of $K$ satisfy $P$. (Note that $K$ may in fact be infinite!)
@@ 2.1. Pointfree topology
As mentioned before, compactness has a very nice interpretation in terms of universal quantification of verifiable properties. We will explore this topic further in section 4.
@@ 2.2. The Scott topology and exponentials
The set of opens $\mathcal{O}X$ of a locale $X$ is itself often endowed with the Scott topology. {defi} A subset $V\subseteq\mathcal{O}X$ is Scott-open if it is upward closed and if whenever a join $\bigvee D$ of a directed set $D$ is in $V$, then some $d\in D$ lies in $V$. Recall that a set $D$ is directed if every finite subset of $D$ has an upper bound in $D$. ∎
The Scott topology can actually be defined more generally on dcpos, which are often used to model data types in programming languages (as discussed in subsection 4.3). {defi} A poset which admits joins of all directed subsets is called a dcpo. We will write $\operatorname*{\bigvee{}^{\!\uparrow}_{\hskip-0.60275pt\hskip 2.2604pt}}S$ for the join of a directed set $S$. A morphism of dcpos or a Scott-continuous function is a monotone map which preserves directed suprema. ∎
@@
Recall if $X,B$ are objects in a category that the exponential object $B^{X}$ is equipped with a morphism $\mathrm{ev}\colon B^{X}\times X\to B$ such that for every $f\colon Y\times X\to B$ there is a unique $\overline{f}\colon Y\to B^{X}$ such that the following diagram commutes.
@@
If $X$ is a locally compact locale, then $X$ is exponentiable (see [13]). In particular, cale $\Sigma^{X}$ exists and its frame is given by the frame of opens of the Scott topology on $\mathcal{O}X\cong\mathrm{Hom}(X,\Sigma)$. Furthermore, $\Sigma^{X}$ is itself locally compact (for instance, see [29, Theorem 11.1]). However, if $X$ is not locally compact, the exponential $\Sigma^{X}$ does not exist. If $X$ is a non-locally-compact topological space, we find that the evaluation map $\mathrm{ev}\colon\mathcal{O}X\times X\to\Sigma$, which sends $(f,x)$ to $f(x)$, is not continuous when $\mathcal{O}X$ is given the Scott topology.
@@ 3.1. Presentations and weak exponentials
Recall that $\langle G\mid R\rangle$ is a quotient of the free frame on $G$, which in turn is the frame of opens of $\Sigma^{G}$. The frame quotient $\Sigma^{G}\twoheadrightarrow\langle G\mid R\rangle$ exhibits $X$ as a sublocale of $\Sigma^{G}$. The points of $\Sigma^{G}$ correspond to subsets of $G$ and we can then interpret this inclusion as sending points of $X$ to the set of generating opens in which they lie. The points in the image of this inclusion correspond to the frame homomorphisms $\mathcal{O}(\Sigma^{G})\to\mathcal{O}1$ which respect the relations. The additional points in $\Sigma^{G}$ may be thought of as ‘imaginary points’ in the same sense that the exponential spaces used in (the different approach of) [7, Chapter 10] are described as imaginary exponentials. Just as complex numbers may be used to more easily prove results about real numbers, so will these imaginary points be useful for our study of compactness in section 4.
@@
Observe that $u$ defines an open in $A\times X$. Since $A\times X$ is a subspace of $A\times\Sigma^{G}$, there exists a (not necessarily unique) open $u^{\prime}$ in $A\times\Sigma^{G}$ which restricts to give $u$. By the universal property of the exponential $\Sigma^{\Sigma^{G}}$, we obtain a morphism $\overline{u}^{\prime}$ making the bottom right triangle commute. The left-hand trapezium commutes since both composites give $\overline{u}^{\prime}\times i_{X}$ and then a simple diagram chase confirms that $u=\widetilde{\mathrm{ev}}\circ(\overline{u}^{\prime}\times X)$.
The above construction is summarised in the following definition. {defi} Let $X$ be a locale with $\mathcal{O}X=\langle G\mid R\rangle$. We define the associated (idealised) machine space to be $\Sigma^{\Sigma^{G}}$ together with the evaluation map $\widetilde{\mathrm{ev}}\colon\Sigma^{\Sigma^{G}}\times X\to\Sigma$ defined in 1. ∎ We explore the intuition behind the concept of machine space in the following section.
@@ 3.2. Philosophical Interlude
Given a space $X$ with $\mathcal{O}X=\langle G\mid R\rangle$ we interpret each generator $g\in G$ as a basic black-box machine which takes as input the points of $X$ and either halts after some finite amount of time, or runs forever. More complicated machines can be constructed from these basic machines. For instance, we may construct a composite machine by taking a finite collection of machines and insist that each machine halts on the given point before the composite machine is considered to have halted. We can also construct a further machine by running any number of these composite machines in parallel and halting on a point if at least one branch halts. This corresponds to taking formal joins of finite formal meets of generators and so we see that these machines are precisely the elements of the free frame on $G$ – that is, the points of the machine space $\Sigma^{\Sigma^{G}}$.
{defi}
@@
Notice that $\Sigma^{\Sigma^{G}}$ by itself contains no information about the behaviour of the machines on inputs from the locale $X$ with presentation $\langle G\mid R\rangle$, as this data is contained in the relation $R$. We can recover this information from the evaluation map $\widetilde{\mathrm{ev}}\colon\Sigma^{\Sigma^{G}}\times X\to\Sigma$ of 1, which records whether a machine halts on a given point.
Remark 2.
In fact, for many purposes any weak exponential with base $\Sigma$ and exponent $X$ behaves like a machine space for $X$. This could be interesting to explore further. However, for this paper we restrict our consideration to $\Sigma^{\Sigma^{G}}$, since the compactness algorithm given in section 4 is dependent on the presentation of the locale by generators and relations.
A point $x\in X$ is accepted by a machine $m$ if there exists some $J_{k}$ such that $x$ is accepted by each ‘basic machine’ $g_{j}$ for all $j\in J_{k}$ (i.e. such that $x$ lies in the open given by $g_{j}$ for all such $j$). In this way, each machine $m$ will define an open on which it halts. Many of these machines will correspond to the same opens, as described by the relation $R$, so that each open in $X$ corresponds to an equivalence class of machines. In subsection 3.3 of subsection 3.3 we will see that when the exponential $\Sigma^{X}$ exists there is a quotient map $\Sigma^{\Sigma^{G}}\twoheadrightarrow\Sigma^{X}$ assigning machines to opens.
Remark 3.
@@
Recall that the failure of the exponential $\Sigma^{X}$ to always exist poses a puzzle for the verifiable interpretation of topology. Luckily, we have seen that machine spaces always exist. It is now helpful to consider in detail how such spaces might correspond to concrete objects that can be learnt about via physical measurements. Note that with such “real-world” models the set of generators is assumed to be countable.
Model 4.
@@
Note that associativity and distributivity are automatically satisfied with distributivity following from the fact that the formal operations are already reduced to the canonical form of a join of meets. These axioms together are then enough to give a frame. ∎
@@
We can now answer the question posed in the introduction. While we have seen that machine space is reasonably concrete, the ‘space of opens’ is less so. In the flawed argument for the existence of a space of opens from section 1, one assumes that given an open and a point, it is possible to semi-decide membership of the point in the open. But an open merely describes the verifiable property under consideration — it is a machine that actually carries out this verification.
In the following subsection, we will show that the quotient map $\Sigma^{\Sigma^{G}}\twoheadrightarrow\mathcal{O}X$ has a section whenever the space $X$ is locally compact. From the above perspective, this fact perfectly explains why in this case a continuous evaluation map exists for $\mathcal{O}X$: given an open $u$ and a point $x$, one can apply the section to $u$ to acquire a machine $s(u)$ which can then verify membership of $x$ in $u$.
@@
Remark 5.
Note that since we do not care which specific machine representing a given open is selected, we can relax the requirement of having an actual section to having a sufficiently well-behaved quotient. In particular, suplattice and preframe homomorphisms between frames are known to correspond to angelically and demonically nondeterministic maps respectively [24, 31, 27] and it would seem that a nondeterministic choice of representative for each open would be sufficient to construct the evaluation map. Indeed, open and proper quotients, or more generally, triquotients, have such ‘non-deterministic sections’ and these are strong enough conditions on the quotient map to entail that $X$ is exponentiable. We will discuss the role of nondeterminism in more detail in a later paper.
@@ 3.3. Relationship between machine space and the space of opens
$\Sigma^{\Sigma^{G}}\times X$$\Sigma^{X}\times X$$\Sigma$$\Sigma^{X}$$\Sigma^{\Sigma^{G}}$$\Sigma^{\Sigma^{G}}\times\Sigma^{G}$$q\times X$$\mathrm{ev}$$\widetilde{\mathrm{ev}}$$\mathrm{ev}$$\Sigma^{\Sigma^{G}}\times i_{X}$$\widetilde{\mathrm{ev}}$$q$
@@
Proposition 6.
Let $X$ be a locally compact locale with $\mathcal{O}X=\langle G\mid R\rangle$. The map $q\colon\Sigma^{\Sigma^{G}}\to\Sigma^{X}$ as defined in subsection 3.3 has a section $s$ satisfying $\widetilde{\mathrm{ev}}\circ(s\times X)=\mathrm{ev}$.
Proof 3.2.
@@
By the diagram we have $\mathrm{ev}\circ(q\times X)\circ(s\times X)=\widetilde{\mathrm{ev}}\circ(s\times X)=\mathrm{ev}=\mathrm{ev}\circ(\mathrm{id}_{\Sigma^{X}}\times X)$. Hence by the uniqueness condition of the universal property of the ‘true’ exponential $\Sigma^{X}$, we have $q\circ s=\mathrm{id}_{\Sigma^{X}}$, as required.
@@
Recall that if a locale $X$ is given by $\mathcal{O}X=\langle G\mid R\rangle$ then the frame $\mathcal{O}X$ is a quotient of $\mathcal{O}(\Sigma^{G})$ by the congruence generated from the relations $R$. Let us call this frame quotient map $\overline{q}$ for notational convenience, but as a function it is the same as $i_{X}^{*}$. Recall that $\mathcal{O}(\Sigma^{G})$ equipped with the Scott topology is precisely the machine space $\Sigma^{\Sigma^{G}}$. Below we will write $\Sigma^{\Sigma^{G}}$ for $\mathcal{O}(\Sigma^{G})$ equipped with the Scott topology even when viewing it as a topological space (as opposed to a locale). There is no danger in doing this since $\Sigma^{\Sigma^{G}}$ is spatial. We can now use $\overline{q}$ to induce a topology on $\mathcal{O}X$. We will make use of the following result.
Lemma 7.
Let $L$ and $M$ be posets admitting all joins and let $q\colon L\twoheadrightarrow M$ be a join-preserving surjection. Then if $L$ is equipped with the Scott-topology, the quotient topology induced on $M$ is also the Scott-topology.
Proof 3.3.
We need to show that $U\subseteq M$ is Scott-open if and only if $q^{-1}(U)$ is Scott-open. If $U$ is Scott-open, then $q^{-1}(U)$ is Scott-open, because $q$ is Scott-continuous (as it preserves all joins).
Now suppose $q^{-1}(U)$ is Scott-open. Consider $a\in U$ and suppose $a\leq b$. Take $a^{\prime}\in q^{-1}(\{a\})$ and $b^{\prime}\in q^{-1}(\{b\})$. We may assume $b^{\prime}\geq a^{\prime}$ by joining $a^{\prime}$ if necessary. Then $a^{\prime}\in q^{-1}(U)$ and hence $b^{\prime}\in q^{-1}(U)$, since $q^{-1}(U)$ is upward closed. Thus, $b=q(b^{\prime})\in U$ and so $U$ is an upset. Next suppose $\operatorname*{\bigvee{}^{\!\uparrow}_{\hskip-0.60275pt\hskip 2.2604pt}}D\in U$. Set $D^{\prime}=q^{-1}({\downarrow}D)$. We note $D^{\prime}$ is directed. Certainly $D^{\prime}$ is nonempty. If $c,d\in D^{\prime}$ then there is an $e\in D$ with $q(c),q(d)\leq e$ by directedness of $D$. Thus, $q(c\vee d)=q(c)\vee q(d)\leq e\in{\downarrow}D$ and so $c\vee d\in D^{\prime}$. Furthermore, note that $q(D^{\prime})={\downarrow}D$ and $q(\operatorname*{\bigvee{}^{\!\uparrow}_{\hskip-0.60275pt\hskip 2.2604pt}}D^{\prime})=\operatorname*{\bigvee{}^{\!\uparrow}_{\hskip-0.60275pt\hskip 2.2604pt}}q(D^{\prime})=\operatorname*{\bigvee{}^{\!\uparrow}_{\hskip-0.60275pt\hskip 2.2604pt}}{\downarrow}D=\operatorname*{\bigvee{}^{\!\uparrow}_{\hskip-0.60275pt\hskip 2.2604pt}}D\in U$. Thus, $\operatorname*{\bigvee{}^{\!\uparrow}_{\hskip-0.60275pt\hskip 2.2604pt}}D^{\prime}\in q^{-1}(U)$ and so there is a $d^{\prime}\in D^{\prime}$ with $d^{\prime}\in q^{-1}(U)$, since $q^{-1}(U)$ is Scott-open. Now $q(d^{\prime})\in{\downarrow}D$ so that $q(d^{\prime})\leq d$ for some $d\in D$. Recalling that $U$ is an upset, we have $d\geq q(d^{\prime})\in U$ and so $d$ is the desired $d\in U$. Thus, $U$ is Scott-open.
Corollary 8.
@@
Lemma 9.
The map $\overline{q}\colon\Sigma^{\Sigma^{G}}\to\mathcal{O}X$ is equal to $q\colon\Sigma^{\Sigma^{G}}\to\Sigma^{X}$ as defined in subsection 3.3 whenever $X$ is locally compact.
Proof 3.4.
@@
The map $\mathrm{Hom}(i_{X},\Sigma)$ corresponds to the action of the frame map $i_{X}^{*}=\overline{q}$, while $\mathrm{Hom}(1,\Sigma^{i_{X}})$ gives the action of $\Sigma^{i_{X}}$ on points. Thus, we have that $\Sigma^{i_{X}}$ and $\overline{q}$ agree on points and hence they coincide.
The failure of $\mathcal{O}X$ with the Scott topology to be an exponential $\Sigma^{X}$ when $X$ is not locally compact is due to this quotient map being badly behaved in general. Indeed, if there were a section to $q$ as in 6 we could show $X$ to be locally compact after all.
Proposition 10.
@@
Proof 3.5.
Since $\Sigma^{G}$ is locally compact, $\mathcal{O}(\Sigma^{G})$ is a continuous frame and in particular a continuous dcpo. But continuous dcpos are known to be stable under Scott-continuous retracts and so $\mathcal{O}X$ is also a continuous dcpo (see [1, Theorem 3.1.4.]). Thus, $\mathcal{O}X$ is a continuous frame and $X$ is locally compact.
To see that $\widetilde{\mathrm{ev}}\circ(s\times X)$ is the evaluation map, first note that, by 9, $\overline{q}=q$ as defined in subsection 3.3. So by the commutative diagram in subsection 3.3, the evaluation map $\mathrm{ev}\colon\mathcal{O}X\times X\to\Sigma$ satisfies $\mathrm{ev}\circ(q\times X)=\widetilde{\mathrm{ev}}$. Hence, $\widetilde{\mathrm{ev}}\circ(s\times X)=\mathrm{ev}\circ(q\times X)\circ(s\times X)=\mathrm{ev}$.
In fact, we can say more (see 5). It is known that the lack of exponential objects is related to the fact that quotients are not stable under products (see [6]). We have the following result. (This result requires us to assume $X$ is spatial and that we are in the category of topological spaces, though it is still useful for intuition. But see also 12.)
Proposition 11.
@@
If $h\colon A\times X\to\Sigma$ then by the weak universal property of $\widetilde{\mathrm{ev}}$ we have a map $h^{\prime}\colon A\to\Sigma^{\Sigma^{G}}$ making the diagram commute. Then $h^{\prime\prime}=\overline{q}h^{\prime}$ satisfies the condition needed for the weak universal property of $\mathrm{ev}\colon\mathcal{O}X\times X\to\Sigma$ to hold. Moreover, it is unique since we have uniqueness on the underlying sets by the universal property of the exponential in $\mathrm{Set}$. Thus, $\mathcal{O}X$ satisfies the universal property of the exponential $\Sigma^{X}$, as required.
Remark 12.